# arXiv:2603.03264v1[astro-ph.SR] 03 Mar 2026

### 03 Mar 2026 v1

@@ Appendix A Data tables
Tables 2, 3 and 4 report our $\delta\langle B\rangle$ and $\delta T$ measurements obtained from spectra recorded with SPIRou for EV Lac, DS Leo and Barnard’s star, respectively. Tables 5 and 6 present similar data obtained from spectra recorded with Narval/ESPaDOnS for EV Lac and DS Leo, respectively.
@@
Table 3: Same as Table 2 for DS Leo.
@@
Table 4: Same as Table 2 for Barnard’s star.
@@
Table 5: $\delta\langle B\rangle$ and $\delta T$ measurements obtained for EV Lac from spectra recorded with Narval/ESPaDOnS. [^4] The complete table is available online in machine readable format.
@@
Table 6: Same as Table 5 for DS Leo.
@@
Appendix B Gaussian Process fit parameters
@@
Table 7: GP hyper-parameters obtained for EV Lac, DS Leo and Barnard’s star from SPIRou spectra. [^5] The hyper-parameters of the kernel are the mean value ($\mu$), the standard deviation of the added uncorrelated noise ($\sigma$), the amplidute ($\alpha$), the decay time ($l$), the smoothing factor $\beta$ and the recurrence period ($P_{\rm rot}$).
@@ Appendix C Additional comparisons
Figure 8 presents the comparison between $\delta\langle B\rangle$ obtained in this paper, and those of Cristofari et al. (2025a). Figure 9 presents comparisons between the retrieved $\delta T$ and the model-driven variations obtained with the LBL.
Figure 8: Same as Fig. 3 for DS Leo (top panel) and Barnard’s star (bottom panel).
Figure 9: Comparison between our model-driven $\delta T$ and the data-driven estimates obtained with the LBL (see Artigau et al. 2024; Cristofari et al. 2025a) for EV Lac (top panel), DS Leo (middle panel) and Barnard’s star (bottom panel). The red line shows the equality.
@@ Appendix D $\delta\langle B\rangle$$\delta T$ relations
Figure 10: Same as Fig. 6 for DS Leo (top panel) and Barnard’s star (bottom panel).
@@ Appendix E Corner plots
Figures 12, 13 and 14 present the corner plots showing the posterior distributions of the hyper-parameters associated with the quasi-periodic Gaussian Fit on $\delta\langle B\rangle$ estimates obtained from SPIRou spectra for EV Lac, DS Leo and Barnard’s star, respectively. Figures 15, 16 present similar corner plots obtained with $\delta\langle B\rangle$ estimates derived from Narval/ESPaDOnS spectra for EV Lac and DS Leo, respectively.
@@
Figure 13: Same as Fig. 12 for DS Leo.
Figure 14: Same as Fig. 12 for Barnard’s star.
Figure 15: Same as Fig. 12 for EV Lac from data recorded with Narval/ESPaDOnS.
Figure 16: Same as Fig. 12 for DS Leo from data recorded with Narval/ESPaDOnS.
@@ Appendix F S-index calculation
Figure 17: Example of spectrum obtained with Narval, showing the V and R rectangular bands, along withe the K and H triangular bands used to compute the S-index.

### 26 Mar 2026 v2

@@ Appendix A Data tables
Tables 2, 3, and 4 report our $\delta\langle B\rangle$ and $\delta T$ measurements obtained from spectra recorded with SPIRou for EV Lac, DS Leo and Barnard’s star, respectively. Tables 5 and 6 present similar data obtained from spectra recorded with Narval/ESPaDOnS for EV Lac and DS Leo, respectively.
@@
Table 3: Same as Table 2 but for DS Leo.
@@
Table 4: Same as Table 2, but for Barnard’s star.
@@
Table 5: $\delta\langle B\rangle$ and $\delta T$ measurements obtained for EV Lac from spectra recorded with Narval/ESPaDOnS (extract). [^4] The complete table is available at the CDS.
@@
Table 6: Same as Table 5 but for DS Leo.
@@
Appendix B GP fit parameters
@@
Table 7: GP hyper-parameters obtained for EV Lac, DS Leo, and Barnard’s star from SPIRou spectra. [^5] The hyper-parameters of the kernel are the mean value ($\mu$), the standard deviation of the added uncorrelated noise ($\sigma$), the amplidute ($\alpha$), the decay time ($l$), the smoothing factor $\beta$ and the recurrence period ($P_{\rm rot}$).
@@ Appendix C Additional comparisons
Figure 8 presents the comparison between $\delta\langle B\rangle$ obtained in this paper, and those of Cristofari et al. (2025). Figure 9 presents comparisons between the retrieved $\delta T$ and the model-driven variations obtained with the LBL.
Figure 8: Same as Fig. 3, but for DS Leo (top panel) and Barnard’s star (bottom panel).
Figure 9: Comparison between our model-driven $\delta T$ and the data-driven estimates obtained with the LBL (see Artigau et al. 2024; Cristofari et al. 2025) for EV Lac (top panel), DS Leo (middle panel), and Barnard’s star (bottom panel). The red line shows the equality.
@@ Appendix D $\delta\langle B\rangle$$\delta T$ relations
Figure 10: Same as Fig. 6 but for DS Leo (top panel) and Barnard’s star (bottom panel).
@@ Appendix E Corner plots
Figures 12, 13, and 14 present the corner plots showing the posterior distributions of the hyper-parameters associated with the quasi-periodic Gaussian Fit on $\delta\langle B\rangle$ estimates obtained from SPIRou spectra for EV Lac, DS Leo and Barnard’s star, respectively. Figures 15 and 16 present similar corner plots obtained with $\delta\langle B\rangle$ estimates derived from Narval/ESPaDOnS spectra for EV Lac and DS Leo, respectively.
@@
Figure 13: Same as Fig. 12, but for DS Leo.
Figure 14: Same as Fig. 12, but for Barnard’s star.
Figure 15: Same as Fig. 12, but for EV Lac from data recorded with Narval/ESPaDOnS.
Figure 16: Same as Fig. 12, but for DS Leo from data recorded with Narval/ESPaDOnS.
@@ Appendix F S-index calculation
Figure 17: Example of spectrum obtained with Narval, showing the V and R rectangular bands, along with the K and H triangular bands used to compute the S-index.